### v1

@@ 2.2 Estimation and Inference of the Minimum Spanning Set
which represents the bootstrap version of $M_{i,T}$. We use the $(1-p)$-quantile of $\max_{i\leq d}M_{i,T}^{b}$, denoted as $\overline{cv}_{1-p,T}^{b}$, as the critical value to construct the one-step estimator of the $\mathrm{MSS}$:
$\overline{\mathrm{MSS}}_{p}\equiv\left\{i=1,\ldots,d\text{: }M_{i,T}>\overline{cv}_{1-p,T}^{b}\right\}.$ (2.15)
Since the $\mathrm{MSS}$ is a subset of $\mathcal{M}$, the critical value $\overline{cv}_{1-p,T}^{b}$ provides an upper bound for the infeasible critical value $cv_{1-p,T}^{u}$. This ensures that the probability of $\overline{\mathrm{MSS}}_{p}$ overestimating the $\mathrm{MSS}$ is less than $p$. The properties of the one-step estimator $\overline{\mathrm{MSS}}_{p}$ are formally presented in the theorem below.
@@
Then we have $\lim_{T\rightarrow\infty}\mathrm{P}(\mathrm{MSS}\subset\overline{\mathrm{MSS}}_{p})=1$ and $\lim\inf_{T\rightarrow\infty}\mathrm{P}(\mathrm{MSS}=\overline{\mathrm{MSS}}_{p})\geq 1-p$.
Theorem 1 shows that the one-step estimator $\overline{\mathrm{MSS}}_{p}$ will not underestimate the $\mathrm{MSS}$, meaning that all assets in the $\mathrm{MSS}$ are included in $\overline{\mathrm{MSS}}_{p}$ wpa1. Additionally, as the sample size increases, $\overline{\mathrm{MSS}}_{p}$ identifies the $\mathrm{MSS}$ with a probability greater than $1-p$. Therefore, the probability that $\overline{\mathrm{MSS}}_{p}$ overestimates the $\mathrm{MSS}$ is asymptotically less than $p$.
In practice, one can choose a small value for $p$, such as 0.05 or 0.01, to ensure that $\overline{\mathrm{MSS}}_{p}$ asymptotically identifies the $\mathrm{MSS}$ with a high probability, such as 0.95 or 0.99. Alternatively, one may set $p=\gamma_{T}$, where $\gamma_{T}$ approaches zero as the sample size increases, ensuring that $\overline{\mathrm{MSS}}_{p}$ identifies the $\mathrm{MSS}$ consistently. Indeed, as we will show below, this is the approach we adopt in the first step of our two-step estimation procedure of the $\mathrm{MSS}$.
The condition in (2.16) specifies the least favorable scenario under which the one-step estimator $\overline{\mathrm{MSS}}_{p}$ will still cover the $\mathrm{MSS}$ wpa1 and overestimates it with probability asymptotically less than $p$. This condition is derived from the identification conditions in (2.2), and is similarly adjusted to resemble the local power analysis of statistical hypothesis test. If there exists a constant $K>0$ such that (2.2) holds for any asset $R_{i}$ in $\mathrm{MSS}$ with
@@
then (2.16) is trivially satisfied. More importantly, (2.16) allows for the identification conditions (2.2) to nearly hold for some assets, in the sense that their values of $\alpha_{i}^{2}+(1_{d-1}^{\top}\beta_{i}-1)^{2}$ are close to zero.
Compared to the desirable estimator $\widehat{\mathrm{MSS}}_{p}^{u}$, which is based on the infeasible critical value $cv_{1-p}^{u}$, it is evident that $\overline{\mathrm{MSS}}_{p}\subset\widehat{\mathrm{MSS}}_{p}^{u}$ with high probability because $cv_{1-p}^{u}\leq\overline{cv}_{1-p,T}^{b}$ wpa1. This implies that $\overline{\mathrm{MSS}}_{p}$ has a higher probability of missing assets in the $\mathrm{MSS}$ compared to $\widehat{\mathrm{MSS}}_{p}^{u}$ in finite samples, although asymptotically, both should include the $\mathrm{MSS}$ wpa1. The reason for this unfavorable property of the one-step estimator $\overline{\mathrm{MSS}}_{p}$ is that its critical value $\overline{cv}_{1-p,T}^{b}$ is asymptotically larger than the value $cv_{1-p}^{u}$ required for exact control of the type-I error (i.e., the error of overestimating the $\mathrm{MSS}$) in statistical inference. This leads to potential power loss in detecting assets in the $\mathrm{MSS}$, especially when the size of the $\mathrm{MSS}$ is relatively large.
To mitigate the power loss of the one-step estimator $\overline{\mathrm{MSS}}_{p}$, we now introduce the two-step procedure for estimating the $\mathrm{MSS}$. In the first step, we obtain a preliminary estimator $\overline{\mathrm{MSS}}_{\gamma_{T}}$ for some positive sequence $\gamma_{T}=o(1)$. [^3] We recommend $\gamma_{T}=\min\{p/2,$ $0.05/\log(\log(T))\}$, which is used in the simulation studies and empirical application of this paper. In view of Theorem 1 with $p=\gamma_{T}$, the preliminary estimator $\overline{\mathrm{MSS}}_{\gamma_{T}}$ should be identical to the $\mathrm{MSS}$ wpa1. This motivates us to estimate the infeasible critical value $cv_{1-p}^{u}$, the $(1-p)$-quantile of $\max_{i\notin\mathrm{MSS}}\tilde{M}_{i,T}$, by the $(1-p)$-quantile of $\max_{i\notin\overline{\mathrm{MSS}}_{\gamma_{T}}}M_{i,T}^{b}$. The latter is denoted as $cv_{1-p,T}^{b}$. When $\overline{\mathrm{MSS}}_{\gamma_{T}}=\mathcal{M}$, we set $cv_{1-p,T}^{b}=0$. Subsequently, our estimate of the $\mathrm{MSS}$ is:
$\widehat{\mathrm{MSS}}_{p}\equiv\left\{i=1,\ldots,d\text{: }M_{i,T}>cv_{1-p,T}^{b}\right\}.$ (2.17)
The implementation details of the two-step procedure are provided in the algorithm below.
@@
Here, $(T^{-1/2}\hat{s}_{i,1}cv_{1-p,T}^{b})_{i\leq d}$ and $(T^{-1/2}\hat{s}_{i,2}cv_{1-p,T}^{b})_{i\leq d}$ can be interpreted as uniform confidence bands for $\left|\alpha_{i}\right|$ and $|1_{d-1}^{\top}\beta_{i}-1|$, respectively. From the expression in (2.18), it follows that an asset $i$ is selected into $\widehat{\mathrm{MSS}}_{p}$, if either $\left|\hat{\alpha}_{i}\right|$, $|1_{d-1}^{\top}\hat{\beta}_{i}-1|$, or both, exceed their respective uniform confidence bands. [^4] As noted in Kan and Zhou (2012), these two cases, i.e., $|\alpha_{i}|>0$ and $|1_{d-1}^{\top}\beta_{i}-1|$, are subject to explicit economic interpretation. Specifically, $|\alpha_{i}|>0$ implies asset $i$ contributes to the tangency portfolio, while $|1_{d-1}^{\top}\beta_{i}-1|>0$ corresponds to the scenario where asset $i$ contributes to the global minimum variance portfolio. For more details on the derivations, see Section 2.1 in Kan and Zhou (2012). Therefore, our method provides detailed insights into the sources of contribution for the selected assets. Additionally, by ranking all assets in the full set according to their $M_{i,T}$ values, we can directly measure their relative importance. [^5] For an empirical illustration of these ideas, refer to Figures 2 and 3, , as well as the related discussion in Section 4.
Algorithm: A Bootstrap $\mathbf{MSS}$ Estimation Procedure
Step 3. Let $m\equiv[T/\ell]$ and $T_{B}\equiv m\ell$. Resample $m$ blocks $Z_{j}^{b}\equiv\{R_{\ell(j-1)+1}^{b},\ldots,R_{\ell(j-1)+\ell}^{b}\}$ independently from $\{B_{j}\}_{j\leq q}$ to obtain a bootstrap sample $\{R_{t}^{b}\}_{t\leq T_{B}}\equiv\{Z_{j}^{b}\}_{j\leq m}$, and then calculate $M_{i,T}^{b}$ specified in (2.14).
Step 4. Repeat Step 3 $B$ times, and set $\overline{cv}_{1-\gamma_{T},T}^{b}$ to the $(1-\gamma_{T})$-sample quantile of $\{\max_{i\leq d}M_{i,T}^{b}\}_{b\leq B}$.
Step 5. Find the preliminary $\mathrm{MSS}$ estimator $\overline{\mathrm{MSS}}_{\gamma_{T}}$ using $\overline{cv}_{1-\gamma_{T},T}^{b}$ and the formula in (2.15).
Step 6. If $\overline{\mathrm{MSS}}_{\gamma_{T}}=\mathcal{M}$, set $cv_{1-p,T}^{b}=0$. Otherwise set $cv_{1-p,T}^{b}$ to the $(1-p)$-sample quantile of $\{\max_{i\notin\overline{\mathrm{MSS}}_{\gamma_{T}}}M_{i,T}^{b}\}_{b\leq B}$, and obtain the $\mathrm{MSS}$ estimator $\widehat{\mathrm{MSS}}_{p}$ using the formula in (2.17).
Step 7. If $\widehat{\mathrm{MSS}}_{p}$ is an empty set in Step 6, set $\widehat{\mathrm{MSS}}_{p}=\{i=1,\ldots,d:M_{i,T}\geq\max_{j\leq d}M_{j,T}\}$.
The last step in the algorithm above is necessary as a finite sample adjustment. When the size of the $\mathrm{MSS}$ is small and the identification condition in (2.2) nearly holds, the test based on the estimated critical value $cv_{1-p,T}^{b}$ may cause the estimator $\widehat{\mathrm{MSS}}_{p}$ to be an empty set with a small sample size. However, since the $\mathrm{MSS}$ cannot be empty, we estimate it as the set of assets that have the highest values of $M_{i,T}$ in this case.
The above algorithm can be simplified to obtain an one-step estimator $\overline{\mathrm{MSS}}_{p}$ that satisfies Theorem 1. Specifically, in Steps 4 and 5 of the algorithm, we replace $\gamma_{T}$ with $p$ to obtain $\overline{\mathrm{MSS}}_{p}$. If $\overline{\mathrm{MSS}}_{p}$ results in an empty set, we redefine it according to Step 7 of the algorithm.
Theorem 2.
Suppose that Assumptions  1 and 2 in the Appendix holds, and moreover
@@
$\min_{i\in\mathrm{MSS}}\left(\alpha_{i}^{2}+\frac{(1_{d-1}^{\top}\beta_{i}-1)^{2}}{d-1}\right)\succ\gamma_{T}^{-1}T^{-1}.$ (2.19)
Then we have $\lim_{T\rightarrow\infty}\mathrm{P}(\mathrm{MSS}\subset\widehat{\mathrm{MSS}}_{p})=1$. Moreover, if $\mathrm{MSS}$ is a proper subset of $\mathcal{M}$, then $\lim_{T\rightarrow\infty}\mathrm{P}(\mathrm{MSS}=\widehat{\mathrm{MSS}}_{p})=1-p$.
Theorem 2 establishes that the two-step $\mathrm{MSS}$ estimator $\widehat{\mathrm{MSS}}_{p}$ possesses the same desirable properties as the infeasible estimator $\widehat{\mathrm{MSS}}_{p}^{u}$, as conjectured in (2.11) and (2.12). Specifically, $\widehat{\mathrm{MSS}}_{p}$ covers the $\mathrm{MSS}$ wpa1, and overestimates the $\mathrm{MSS}$ with probability exactly $p$ as the sample size $T$ increases. If $\mathrm{MSS}=\mathcal{M}$, then the first result of Theorem 2 implies that $\lim_{T\rightarrow\infty}\mathrm{P}(\mathrm{MSS}=\widehat{\mathrm{MSS}}_{p})=1$.
Theorem 2 also shows that the two-step $\mathrm{MSS}$ estimator $\widehat{\mathrm{MSS}}_{p}$ shares similar properties with $\widehat{\mathrm{MSS}}_{p}^{u}$ when compared to the one-step estimator $\overline{\mathrm{MSS}}_{p}$. Since $cv_{1-p,T}^{b}\leq\overline{cv}_{1-p,T}^{b}$ by construction, the one-step estimator $\overline{\mathrm{MSS}}_{p}$ is always a subset of the two-step estimator $\widehat{\mathrm{MSS}}_{p}$ in finite samples, although both are expected to cover the $\mathrm{MSS}$ wpa1. This implies that, similar to the infeasible estimator $\widehat{\mathrm{MSS}}_{p}^{u}$, the probability that $\widehat{\mathrm{MSS}}_{p}$ misses any assets in the $\mathrm{MSS}$ is smaller than that of $\overline{\mathrm{MSS}}_{p}$.

### v3

@@ 2.2 Estimation and Inference of the Minimum Spanning Set
where $\hat{\alpha}_{i}^{b}$, $\hat{\beta}_{i}^{b}$ and $\hat{s}_{i,j}^{b}$ denote the bootstrap counterparts of $\hat{\alpha}_{i}$, $\hat{\beta}_{i}$ and $\hat{s}_{i,j}$, respectively.
Step 4. Repeat Step 3 $B$ times to obtain $\{(M_{i,T}^{b})_{i\leq d}\}_{b\leq B}$.
Step 5. For any $\mathcal{S}\subseteq\mathcal{M}_{d}$, let $cv_{1-p,T}^{b}(\mathcal{S})$ denote the $(1-p)$-conditional quantile of $\max_{i\in\mathcal{S}}M_{i,T}^{b}$ given the data. Starting with $\widehat{\mathrm{MSS}}_{p,0}=\varnothing$, we update the estimate of the $\mathrm{MSS}$ through:
$\widehat{\mathrm{MSS}}_{p,j}=\left\{i\in\mathcal{M}_{d}\text{: }M_{i,T}>cv_{1-p,T}^{b}(\mathcal{M}_{d}\backslash\widehat{\mathrm{MSS}}_{p,j-1})\right\}.$ (2.13)
The updating process stops at $j^{\ast}$ when $\widehat{\mathrm{MSS}}_{p,j^{\ast}}=\widehat{\mathrm{MSS}}_{p,j^{\ast}-1}$ or $\widehat{\mathrm{MSS}}_{p,j^{\ast}}=\mathcal{M}_{d}$.
Step 6. The final $\mathrm{MSS}$ estimate is then defined as $\widehat{\mathrm{MSS}}_{p}=\widehat{\mathrm{MSS}}_{p,j^{\ast}}$.
Steps 2–4 in the algorithm above describe the MBB procedure for obtaining the bootstrap statistics $(M_{i,T}^{b})_{i\leq d}$, while Step 5 represents the step-down procedure for constructing the critical value used in our $\mathrm{MSS}$ estimation procedure. Our $\mathrm{MSS}$ estimator can be formally defined as:
$\widehat{\mathrm{MSS}}_{p}\equiv\left\{i\in\mathcal{M}_{d}\text{: }M_{i,T}>cv_{1-p,T}^{b}\right\},$ (2.14)
where $cv_{1-p,T}^{b}\equiv cv_{1-p,T}^{b}(\mathcal{M}_{d}\backslash\widehat{\mathrm{MSS}}_{p,j^{\ast}})$. [^3] We set $cv_{1-p,T}^{b}$ to zero if $\widehat{\mathrm{MSS}}_{p,j^{\ast}}=\mathcal{M}$. When the size of the $\mathrm{MSS}$ is small and the identification condition in (2.2) is nearly satisfied, the test based on $cv_{1-p,T}^{b}$may cause the estimator $\widehat{\mathrm{MSS}}_{p}$ to be an empty set, particularly in small sample sizes. However, since the $\mathrm{MSS}$ cannot be empty, we implement a finite-sample adjustment by estimating it as the set of assets with the highest values of $M_{i,T}$ in such cases. [^4] In the proof of Theorem 1, we show that $\widehat{\mathrm{MSS}}_{p}$, as obtained from our estimation algorithm, is nonempty wpa1. Consequently, the finite-sample adjustment becomes asymptotically negligible.
@@
Then we have $\lim_{T\rightarrow\infty}\mathrm{P}(\mathrm{MSS}\subseteq\widehat{\mathrm{MSS}}_{p})=1$. Moreover, if $\mathrm{MSS}$ is a proper subset of $\mathcal{M}_{d}$, then $\lim_{T\rightarrow\infty}\mathrm{P}(\mathrm{MSS}=\widehat{\mathrm{MSS}}_{p})=1-p$.
Theorem 1 establishes that the $\mathrm{MSS}$ estimator $\widehat{\mathrm{MSS}}_{p}$ retains the desirable properties of covering the true $\mathrm{MSS}$ wpa1, and overestimates the $\mathrm{MSS}$ with probability exactly $p$ as the sample size $T$ increases. If $\mathrm{MSS}=\mathcal{M}_{d}$, then the first result of Theorem 1 implies that $\lim_{T\rightarrow\infty}\mathrm{P}(\mathrm{MSS}=\widehat{\mathrm{MSS}}_{p})=1$. The condition in (2.15) specifies the least favorable scenario under which these desirable properties still hold. This condition is derived from the identification conditions in (2.2), and is similarly adjusted to resemble the local power analysis of statistical hypothesis test. If there exists a constant $K>0$ such that (2.2) holds for any asset $R_{i}$ in $\mathrm{MSS}$ with
@@
then (2.15) is trivially satisfied. More importantly, (2.15) allows for the identification conditions (2.2) to nearly hold, in the sense that some of $\alpha_{i}^{2}+(1_{d-1}^{\top}\beta_{i}-1)^{2}$ ($i\in\mathrm{MSS}$) are close to zero.
@@
Here, [0,[0,[ 0 , $T^{-1/2}\hat{s}_{i,1}cv_{1-p,T}^{b}]$ and [0[0[ 0, $T^{-1/2}\hat{s}_{i,2}cv_{1-p,T}^{b}]$ can be interpreted as uniform confidence bands for estimating the $\mathrm{MSS}$ through testing $\alpha_{i}=0$ and $01_{d-1}^{\top}\beta_{i}-1=$ 0 over $i\in\mathcal{M}_{d}$. From the expression in (2.16), it follows that an asset $i$ is selected into $\widehat{\mathrm{MSS}}_{p}$, if either $\left|\hat{\alpha}_{i}\right|$, $|1_{d-1}^{\top}\hat{\beta}_{i}-1|$, or both, exceed their respective upper bound of the uniform confidence band. [^5] As noted in Kan and Zhou (2012), these two cases, i.e., $|\alpha_{i}|>0$ and $|1_{d-1}^{\top}\beta_{i}-1|>0$, are subject to explicit economic interpretation. Specifically, $|\alpha_{i}|>0$ implies asset $i$ contributes to the tangency portfolio, while $|1_{d-1}^{\top}\beta_{i}-1|>0$ corresponds to the scenario where asset $i$ contributes to the global minimum variance portfolio. For more details on the derivations, see Section 2.1 in Kan and Zhou (2012). Therefore, our method provides detailed insights into the sources of contribution for the selected assets. Additionally, by ranking all assets in the full set according to their $M_{i,T}$ values, we can directly measure their relative importance. [^6] For an empirical illustration of these ideas, refer to Figures 1 and 2, as well as the related discussion in Section 4.
2.3 Other Applications
Although the procedure in the previous section is primarily designed for inference on the $\mathrm{MSS}$, it can be adapted to conduct statistical inference for other unknown parameters of practical interest. In this subsection, we present four illustrative examples.
First, in the proof of Lemma 1 in the Appendix, we show that the $\mathrm{MSS}$ is equivalent to the union of the supports of the tangency portfolio and the global minimum variance portfolio, which correspond to the supports of $(\alpha_{i})_{i\leq d}$ and $(1_{d-1}^{\top}\beta_{i}-1)_{i\leq d}$, respectively. Accordingly, our inference procedure for the $\mathrm{MSS}$ can be slightly modified to perform inference on the minimal asset sets required for constructing these two important portfolios.
We next describe the modifications needed to adapt the $\mathrm{MSS}$ inference procedures for the tangency portfolio support, defined as $\mathrm{TAN}\equiv\{i\in\mathcal{M}_{d}:\alpha_{i}\neq 0\}$. For $i\in\mathcal{M}_{d}$, we define
$M_{i,T}(\alpha)\equiv\frac{T^{1/2}\left|\hat{\alpha}_{i}\right|}{\hat{s}_{i,1}}\text{ \ \ \ and \ \ \ }M_{i,T}^{b}(\alpha)\equiv\frac{T^{1/2}\left|\hat{\alpha}_{i}^{b}-\hat{\alpha}_{i}\right|}{\hat{s}_{i,1}^{b}}.$ (2.17)
By applying the algorithm from the previous section to estimate the $\mathrm{MSS}$, but replacing $M_{i,T}$ and $M_{i,T}^{b}$ by $M_{i,T}(\alpha)$ and $M_{i,T}^{b}(\alpha)$, respectively, we obtain the bootstrap critical value $cv_{1-p,T}^{b}(\alpha)$. This leads to the following estimator for $\mathrm{TAN}$:
$\widehat{\mathrm{TAN}}_{p}\equiv\left\{i\in\mathcal{M}_{d}\text{: }M_{i,T}(\alpha)>cv_{1-p,T}^{b}(\alpha)\right\}.$ (2.18)
Using the arguments presented in the proof of Theorem 1, it can be shown that: $\widehat{\mathrm{TAN}}_{p}$ covers $\mathrm{TAN}$ with probability approaching 1 and is identical to $\mathrm{TAN}$ with probability converging to $1-p$.
The estimation procedure for the support of the global minimum variance portfolio, defined as $\mathrm{GMV}\equiv\{i\in\mathcal{M}_{d}:1_{d-1}^{\top}\beta_{i}\neq 1\}$ is constructed analogously. Specifically, by replacing $M_{i,T}(\alpha)$ and $M_{i,T}^{b}(\alpha)$ in (2.17) with
$M_{i,T}(\beta)\equiv\frac{T^{1/2}\left|1_{d-1}^{\top}\hat{\beta}_{i}-1\right|}{\hat{s}_{i,2}}\text{ \ and \ }M_{i,T}^{b}(\beta)\equiv\frac{T^{1/2}\left|1_{d-1}^{\top}(\hat{\beta}_{i}^{b}-\hat{\beta}_{i})\right|}{\hat{s}_{i,2}^{b}},$ (2.19)
for $i\in\mathcal{M}_{d}$, one can derive the corresponding estimation/inference procedure for $\mathrm{GMV}$. For brevity, we omit the detailed exposition.
In practice, researchers may also be interested in identifying the subset of $\mathrm{TAN}$ consisting of assets with positive $\alpha_{i}$. We denote this set as $\mathrm{TAN}^{+}\equiv\{i\in\mathcal{M}_{d}:\alpha_{i}>0\}$. To obtain estimation/inference procedures for this set, we define
$\tilde{M}_{i,T}(\alpha)\equiv\frac{T^{1/2}\hat{\alpha}_{i}}{\hat{s}_{i,1}}\text{ \ \ \ and \ \ \ }\tilde{M}_{i,T}^{b}(\alpha)\equiv\frac{T^{1/2}(\hat{\alpha}_{i}^{b}-\hat{\alpha}_{i})}{\hat{s}_{i,1}^{b}}.$ (2.20)
Applying the algorithm from the previous section to estimate the $\mathrm{MSS}$, but replacing $M_{i,T}$ and $M_{i,T}^{b}$ by $\tilde{M}_{i,T}(\alpha)$ and $\tilde{M}_{i,T}^{b}(\alpha)$, respectively, we obtain the bootstrap critical value $cv_{1-p,T}^{+,b}(\alpha)$. This leads to the following estimator for $\mathrm{TAN}^{+}$:
$\widehat{\mathrm{TAN}}_{p}^{+}\equiv\left\{i\in\mathcal{M}_{d}\text{: }\tilde{M}_{i,T}(\alpha)>cv_{1-p,T}^{+,b}(\alpha)\right\}.$
Similar results to those stated in Theorem 1 can be established for $\widehat{\mathrm{TAN}}_{p}^{+}$ for valid inference on $\mathrm{TAN}^{+}$.
Finally, our procedure can be modified to perform joint inference on the redundancy of an additional set of assets, denoted as $R_{N}\equiv(R_{i})_{i\in\mathcal{N}}$ relative to a benchmark set of assets, $R_{K}\equiv(R_{i})_{i\in\mathcal{K}}$, where $\mathcal{N}$ and $\mathcal{K}$ form a mutually exclusive and exhaustive partition of $\mathcal{M}_{d}$. For each asset $R_{i}$ in $R_{N}$, we consider the following regression:
$R_{i}=\alpha_{K,i}+\beta_{K,i}^{\top}R_{K}+\varepsilon_{K,i},$
from which we define the test statistic:
@@ 2.3 Other Applications
$M_{i,T}(\mathcal{N})\equiv\max\left\{\frac{T^{1/2}\left|\hat{\alpha}_{K,i}\right|}{\hat{s}_{K,i,1}},\text{ \ }\frac{T^{1/2}\left|1_{k}^{\top}\hat{\beta}_{K,i}-1\right|}{\hat{s}_{K,i,2}}\right\},$ (2.21)
where $\hat{s}_{K,i,j}$ is constructed analogously to $\hat{s}_{i,j}$ ($j=1,2$). To determine the critical value for inference, we define the bootstrap counterpart:
$M_{i,T}^{b}(\mathcal{N})\equiv\max\left\{\frac{T^{1/2}\left|\hat{\alpha}_{K,i}^{b}-\hat{\alpha}_{K,i}\right|}{\hat{s}_{K,i,1}^{b}},\text{ \ }\frac{T^{1/2}\left|1_{k}^{\top}(\hat{\beta}_{K,i}^{b}-\hat{\beta}_{K,i})\right|}{\hat{s}_{K,i,2}^{b}}\right\},$ (2.22)
where $k$ is the cardinality of $\mathcal{K}$, $\hat{\alpha}_{K,i}^{b}$, $\hat{\beta}_{K,i}^{b}$, $\hat{s}_{K,i,1}^{b}$ and $\hat{s}_{K,i,2}^{b}$ are the bootstrap versions of $\hat{\alpha}_{K,i}$, $\hat{\beta}_{K,i}$, $\hat{s}_{K,i,1}$ and $\hat{s}_{K,i,2}$, respectively.
Let $\mathcal{N}^{\ast}$ denote the subset of $\mathcal{N}$ consisting of nonredundant assets, i.e., those that contribute to expanding the mean-variance efficient frontier of $R_{K}$. We can estimate $\mathcal{N}^{\ast}$ by
$\widehat{\mathcal{N}}_{p}^{\ast}\equiv\left\{i\in\mathcal{N}\text{: \ }M_{i,T}(\mathcal{N})>cv_{1-p,T}^{b}(\mathcal{N})\right\},$
where $cv_{1-p,T}^{b}(\mathcal{N})$ denotes the bootstrap critical value from the algorithm in the previous section with $M_{i,T}$ and $M_{i,T}^{b}$ replaced by $M_{i,T}(\mathcal{N})$ and $M_{i,T}^{b}(\mathcal{N})$, respectively. The theoretical properties established for $\mathrm{MSS}$ in Theorem 1 also hold for $\widehat{\mathcal{N}}_{p}^{\ast}$. Specifically, $\widehat{\mathcal{N}}_{p}^{\ast}$ covers $\mathcal{N}^{\ast}$ wpa1, and is identical to $\mathcal{N}^{\ast}$ with probability converging to $1-p$.